### v3

# Conformal e-prediction

Abstract
This paper discusses a counterpart of conformal prediction for e-values, conformal e-prediction. Conformal e-prediction is conceptually simpler and had been developed in the 1990s as precursor of conformal prediction. When conformal prediction emerged as result of replacing e-values by p-values, it seemed to have important advantages over conformal e-prediction without obvious disadvantages. This paper re-examines relations between conformal prediction and conformal e-prediction systematically from a modern perspective. Conformal e-prediction has advantages of its own, such as the ease of designing conditional conformal e-predictors and the guaranteed validity of cross-conformal e-predictors (whereas for cross-conformal predictors validity is only an empirical fact and can be broken with excessive randomization). Even where conformal prediction has clear advantages, conformal e-prediction can often emulate those advantages, more or less successfully. Conformal e-prediction can also serve as basis for testing; the resulting conformal e-testing looks very different from but inherits some strengths of conformal testing.
The version of this paper at http://alrw.net (Working Paper 26) is updated most often.
1 Introduction
Conformal prediction is based on the notion of a p-value. At this time p-values are widely discussed and sometimes criticized (see, e.g., [33]), and several alternatives to p-values have been proposed. Perhaps the most popular alternatives are Bayes factors and their non-Bayesian variation, e-values. The terminology of e-values was introduced in [32], and the literature on e-values has been growing quickly; see, e.g., [18, 17, 7].
In fact, e-values were used (under different names) when discussing precursors of conformal prediction in the 1990s; in this paper we will refer to this variety of conformal prediction as conformal e-prediction. One early description of conformal e-prediction is [5], and [1] is a recent rediscovery. The paper [27] that first introduced conformal prediction also discusses conformal e-prediction. In this paper, we will occasionally refer to conformal prediction as conformal p-prediction in order to emphasize it being based on p-values.
Soon after the publication of [5, 27], conformal e-prediction seems to have disappeared. Perhaps the main reason why it was superseded by conformal prediction was that conformal predictions can be packaged as prediction sets [28, Sect. 2.2], and in this case their property of validity is very easy to state: we just say that the probability of error is at most $\epsilon$ at a prespecified significance level $\epsilon$ [28, Proposition 2.3]. This was clearly stated only in 2001 [12, Theorem 1], although this statement was implicit in the standard requirement of validity for p-values stated in [27]. The standard requirement of validity for e-values, also stated in [27], does not admit such an intuitive restatement in terms of probability of error.
Another reason for conformal e-prediction losing its popularity may have been the finding in 2002 [23, Theorem 1] that, in the on-line mode of prediction, smoothed conformal predictors make errors independently. An important corollary of this stronger property of validity is that small probabilities of errors manifest themselves, with high probability, as a low frequency of errors [28, Corollary 2.5].
A third advantage of conformal prediction is that the strong property of validity found in [23, Theorem 1] allows us to test the assumption of exchangeability. This strong property can be stated as the independence of the conformal p-values output at different steps, and to test exchangeability, we can bet against the conformal p-values being independent and uniformly distributed. This led to the introduction in 2003 [29] of conformal test martingales. For a long time, nothing was known about the efficiency of conformal test martingales, and first results about their efficiency appeared in 2019 (see [26]); this is a major topic of [28, Part III].
The last advantage of conformal prediction that we discuss in this section was found only in 2017 [30], and so it did not contribute to the eclipse of conformal e-prediction in the early 2000s. It was the discovery of conformal predictive distributions, motivated by [19]: in the case of regression, smoothed conformal prediction produces “conformal predictive distributions”, which are automatically well-calibrated.
In this paper we will look systematically at these advantages of conformal prediction except for the last one, which will only be briefly discussed in the concluding section. On one hand, the favourable properties of conformal prediction are often partially satisfied by conformal e-prediction. And on the other hand, conformal e-prediction has several advantages of its own.
For simplicity, in this paper we consider IID (or at least exchangeable) data and concentrate on the problem of pattern recognition (also known as classification). We start in Sect. 2 from the definition of conformal e-prediction and continue in Sect. 3 with discussing its validity. The simplest property of validity (Proposition 1 in Sect. 3) consists in conformal e-predictors producing at each step a valid e-value for the true label: namely, its expectation is at most 1 at each step; the general notion of e-value is introduced right after Proposition 1. Since the expectation is the average over the sample space, we can say that conformal e-predictors are valid in the space domain, or space-wise. A complementary notion of efficiency is efficiency in the time domain, which we consider next.
Proposition 1 does not say anything about the relation between the e-values for the true labels produced at different steps. Is it possible that for some streams of data the average of the e-values produced at different steps tends to 2 while for others it tends to 0.5 with non-zero probability? In Sect. 3 we show that this is impossible in the online prediction protocol, stating both the strong law of large numbers and the law of the iterated logarithm for the e-values produced for the true labels at different steps. This does not fully replace the strong property of independence of errors for conformal prediction, but it can be regarded as a partial replacement. We can see that the e-values produced at different steps are not misleading in the time domain; not only is their expectation at most 1 at each step, their average is at most 1 time-wise in the long run.
The properties of validity established in Sect. 3 are marginal, in the sense that the expectations, or time averages, in them are not conditional on any properties of the observations. For example, if our predictions are for people, in principle we can get very different averages for men and women. In the case of conformal prediction, a simple way of achieving conditional validity is using Mondrian conformal prediction [28, Sect. 4.6]. Mondrian conformal prediction requires a hard partition of observations, such as the partition of people into men and women. Interestingly, conformal e-prediction is much more flexible when trying to achieve conditional validity. Its conditional version does not have to be based on a partition (Sect. 4); e.g., we may require separate validity for men, women, and Europeans. It has been shown recently that this type of conditionality can also be achieved for conformal prediction (see, e.g., [6]), but it is much easier to achieve and appears to be more natural in the case of conformal e-prediction. This can be regarded as an advantage of conformal e-prediction.
What we discuss in Sections 24 is “full” conformal e-prediction, and it is computationally inefficient for many standard ways of producing predictions. Section 5 introduces split conformal e-prediction, which is a simple way to make conformal e-prediction computationally efficient. Similarly to split-conformal prediction (introduced in [15, 14]), split conformal e-prediction can lose in predictive efficiency as compared with full conformal e-prediction.
To prevent loss in computational efficiency without sacrificing predictive efficiency, cross-conformal predictors were introduced in [25]. Cross-conformal predictors are not provably valid [25, Appendix], and this sometimes even shows in experimental results [11]. The limits of violations of validity are given by Rüschendorf’s result (see, e.g., [31]): when merging p-values coming from different folds by taking arithmetic mean (this is essentially what cross-conformal predictors do), the resulting arithmetic mean has to be multiplied by 2 in order to guarantee validity. In the more recent method of jackknife+, introduced in [2] and closely related to cross-conformal prediction, there is a similar factor of 2 [2, Theorem 1], which cannot be removed in general [2, Theorem 2].
In Sect. 6, we introduce a version of cross-conformal prediction based on e-values, which we call cross-conformal e-prediction. The situation with cross-conformal e-prediction is very different from cross-conformal prediction, as the arithmetic mean of e-values is always an e-value. This is an obvious fact, but it is shown in [32, Sect. 3] that arithmetic mean is the only useful merging rule. Therefore, cross-conformal e-prediction is always valid. This is a second advantage of conformal e-prediction.
The emphasis of Sections 36 is on the validity of conformal e-prediction and its computational efficiency, while Sect. 7 moves on to its predictive efficiency. What are suitable criteria of predictive efficiency? We propose two such criteria in the case of pattern recognition, the “observed log criterion” and the “prior log criterion”.
In Sections 8 and 9 we will discuss various ways of testing the exchangeability assumption, one-off and “multistage”, based on conformal e-prediction. We will see that conformal e-testing inherits some strengths of conformal testing.
Section 10 concludes and lists some advantages and disadvantages of conformal e-prediction as compared with conformal prediction.
2 Conformal e-predictors
Suppose we are given a training set $z_{1},\dots,z_{n}$ consisting of labelled objects $z_{i}=(x_{i},y_{i})$ and our goal is to predict the label of a new object $x$. In this paper we consider predictors of the following type: for each potential label $y$ for $x$ we would like to have a number $f(z_{1},\dots,z_{n},x,y)$ reflecting the plausibility of $y$ being the true label of $x$. An example is conformal transducers [28, Sect. 2.7], which, in the terminology of this paper, may be called conformal p-predictors. The output
$y\mapsto f(z_{1},\dots,z_{n},x,y)$
of a conformal p-predictor is the full conformal prediction for the label of $x$; e.g., it determines the prediction set at each significance level. We will sometimes write $f(z_{1},\dots,z_{n},z)$, where $z:=(x,y)$, instead of $f(z_{1},\dots,z_{n},x,y)$.
We will use the notation $\mathbf{X}$ for the object space and $\mathbf{Y}$ for the label space (both assumed non-empty). These are measurable spaces from which the objects and labels, respectively, are drawn. Full observations $z=(x,y)$ are drawn from the observation space $\mathbf{Z}:=\mathbf{X}\times\mathbf{Y}$. For any non-empty set $X$, $X^{+}$ will be the set $\cup_{n=1}^{\infty}X^{n}$ of all non-empty finite sequences of elements of $X$.
A nonconformity e-measure is a measurable function $A:\mathbf{Z}^{+}\to[0,\infty)^{+}$ that maps any finite sequence $(z_{1},\dots,z_{m})$, $m\in\{1,2,\dots\}$, to a finite sequence $(\alpha_{1},\dots,\alpha_{m})$ of the same length consisting of nonnegative numbers with average at most 1,
$\frac{1}{m}\sum_{i=1}^{m}\alpha_{i}\leq 1,$ (1)
that satisfies the following property of equivariance: for any $m\in\{2,3,\dots\}$, any permutation $\pi$ of $\{1,\dots,m\}$, any $(z_{1},\dots,z_{m})\in\mathbf{Z}^{m}$, and any $(\alpha_{1},\dots,\alpha_{m})\in[0,\infty)^{m}$,
$(\alpha_{1},\dots,\alpha_{m})=A(z_{1},\dots,z_{m})\Longrightarrow(\alpha_{\pi(1)},\dots,\alpha_{\pi(m)})=A(z_{\pi(1)},\dots,z_{\pi(m)}).$
The conformal e-predictor $f$ corresponding to such $A$ is defined by
$f(z_{1},\dots,z_{n},x,y):=\alpha_{n+1},\text{\enspace where\enspace}(\alpha_{1},\dots,\alpha_{n},\alpha_{n+1}):=A(z_{1},\dots,z_{n},(x,y)),$ (2)
so that $f:\mathbf{Z}^{+}\to[0,\infty)$. A conformal e-predictor is a function that can be obtained from a nonconformity e-measure in this way.
When given a training set $z_{1},\dots,z_{n}$ and a test object $x$, the full prediction for $x$ according to a conformal e-predictor $f$ is the family of conformal e-values
$\left(f(z_{1},\dots,z_{n},x,y)\mid y\in\mathbf{Y}\right).$
The full prediction for the label of $x$ can be summarized as, e.g., the point prediction
$\hat{y}\in\arg\min_{y}f(z_{1},\dots,z_{n},x,y)$
(assuming the $\min$ is attained at a single label), the e-confidence
$\min_{y\neq\hat{y}}f(z_{1},\dots,z_{n},x,y),$
and the e-credibility $f(z_{1},\dots,z_{n},x,\hat{y})$. We can make a confident point prediction when the e-confidence is large while the e-credibility is not.
Let us say that a nonconformity e-measure and the corresponding conformal e-predictor are admissible if we always have “$=$” in place of $\leq$” in the definition (1). If a conformal e-predictor is not admissible, we can make its predictions more confident without sacrificing their validity. Therefore, we will usually concentrate on admissible nonconformity e-measure and admissible conformal e-predictors.
A nonnegative nonconformity measure $A:\mathbf{Z}^{+}\to[0,\infty)^{+}$ is defined as a nonconformity e-measure except that the condition (1) is omitted. Given a nonnegative nonconformity measure $A$, we can always define the corresponding nonconformity e-measure $A^{\prime}$ by normalizing $A$:
$A^{\prime}(z_{1},\dots,z_{m}):=\frac{m}{\sum_{i=1}^{m}\alpha_{i}}(\alpha_{1},\dots,\alpha_{m}),$ (3)
where $(\alpha_{1},\dots,\alpha_{m}):=A(z_{1},\dots,z_{m})$; if $A(z_{1},\dots,z_{m})=(0,\dots,0)$, we set $A^{\prime}(z_{1},\dots,z_{m}):=(1,\dots,1)$ in order to ensure that $A^{\prime}$ is admissible. We will say that the corresponding conformal e-predictor is based on $A$.
A further generalization of nonnegative nonconformity measures, nonconformity measures, is obtained by dropping the condition of nonnegativity; these are functions $A:\mathbf{Z}^{+}\to\mathbb{R}$. They are used, explicitly or implicitly, in conformal p-prediction.
The conformal e-predictor proposed in [5] for binary pattern recognition problems (with $\mathbf{Y}=\{-1,1\}$) is based on support vector machines (SVM); let us fix all relevant parameters, such as the kernel. It is defined as
$f(z_{1},\dots,z_{n+1}):=\begin{cases}(n+1)/\left|\textrm{SV}\right|&\text{if $n+1\in\textrm{SV}$}\\ 0&\text{otherwise},\end{cases}$
where SV is the set of indices of support vectors: $i\in\textrm{SV}$ if and only if $z_{i}$, $i=1,\dots,n+1$, is a support vector for the SVM constructed from $z_{1},\dots,z_{n+1}$ as training set. It is based on the indicator function of being a support vector. When given a training set $z_{1},\dots,z_{n}$ and a new object $x$, this conformal e-predictor goes through all potential labels $y$ for $x$ and for each constructs an SVM and outputs $f(z_{1},\dots,z_{n},x,y)$. It makes it computationally inefficient.
3 Validity of conformal e-predictors
The following obvious proposition asserts the validity of conformal e-predictors. Let us write $Z_{1},Z_{2},\dots$ for the random vectors whose realizations are the observed data $z_{1},z_{2},\dots$; more generally, $(X,Y)$ or $Z$ are random vectors with values in the observation space $\mathbf{Z}$. As usual, a finite sequence of random vectors is exchangeable if its joint distribution does not change if it is permuted (and an infinite sequence is exchangeable if its joint distribution does not change if its first $n$ elements are permuted, for any $n$). The difference between exchangeability and being IID (independent and identically distributed) can be substantial for finite data sequences (but for infinite data sequences the difference between the two assumptions disappears according to de Finetti’s theorem [28, Sect. A.5.1]).
Proposition 1.
For any conformal e-predictor $f$ and any $n$, if $Z_{1},\dots,Z_{n},(X,Y)$ are IID (or exchangeable),
$\mathbb{E}f(Z_{1},\dots,Z_{n},X,Y)\leq 1$ (4)
(with “$=$” in place of $\leq$” if $f$ is admissible).
Proof.
It follows from the definition of conformal e-predictors that
$\mathbb{E}\left(f(Z_{1},\dots,Z_{n},X,Y)\mid\lbag Z_{1},\dots,Z_{n},(X,Y)\rbag\right)\leq 1,$
and it remains to average over the multisets $\lbag Z_{1},\dots,Z_{n},(X,Y)\rbag$. ∎
The property of validity given in Proposition 1 says that conformal e-predictors output valid e-values for the true labels. Formally, an e-variable is a random variable $E$ satisfying $\mathbb{E}(E)\leq 1$ under the true data-generating distribution (and the values it takes are e-values). We do not expect the e-values for the true labels to be large because, by Markov’s inequality, $\mathbb{P}(E\geq C)\leq 1/C$ for any constant $C>1$.
Proposition 1 involves a “space average”, i.e., an average over the sample space. The analogous property of validity for conformal prediction can be stated in terms of error probabilities for set predictions (see, e.g., [28, Proposition 2.1], which is particularly intuitive. This is the first advantage of conformal prediction pointed out in Sect. 1. It disappears when we move to conformal e-prediction: validity has to be defined in terms of e-values (however, it has been argued [18] that e-values are more intuitive than p-values).
Another advantage of conformal prediction is that, in the online mode of prediction (to be defined shortly), conformal predictors (in their smoothed version) make errors at different steps independently (see, e.g., [28, Theorem 11.1]). Without independence, it is possible, even when the probability of error at each step is $\epsilon$, for the long-term relative frequency of errors over consecutive steps to be either 0 or 1 (1 with probability $\epsilon$). The independence of errors forces the long-term frequency of errors to be $\epsilon$ almost surely (which is also asserted in [28, Proposition 2.1]). We can say that independence implies ergodicity, i.e., the almost sure coincidence of time and space averages.
The independence of conformal p-values in the online mode of prediction is the strongest property of validity in conformal prediction as presented in [28]. In conformal e-prediction, independence is lost, but ergodicity still holds, at least for bounded conformal e-predictors.
To give an example demonstrating the loss of independence in conformal e-prediction, we first define the online prediction protocol. In the online protocol for conformal e-prediction, we observe an object $x_{1}$, apply the conformal e/predictor to compute the conformal e-values $f(x_{1},y)$ for all possible labels $y\in\mathbf{Y}$, observe the true label $y_{1}$, record the e-value $e_{1}:=f(x_{1},y_{1})$ for it, observe another object $x_{2}$, apply the conformal e-predictor to compute the e-values $f(x_{1},y_{1},x_{2},y)$ for all possible labels $y\in\mathbf{Y}$, observe the true label $y_{2}$, record the e-value $e_{2}:=f(x_{1},y_{1},x_{2},y_{2})$ for it, etc. In the case conformal prediction, we can instead record conformal p-values $p_{n}$ (or record whether a mistake is made at a given significance level). While the conformal p-values are independent (for smoothed conformal predictors), the following example shows that they are not already in a toy binary situation (perhaps this is the simplest non-trivial example of conformal e-prediction). We write $E_{n}$ for the e-value $e_{n}$ regarded as a random variable (an e-variable).
Example 2.
Let the random observations $Z_{1},Z_{2},\dots$ correspond to tossing a fair coin: $Z_{n}\in\{0,1\}$ and $Z_{n}=1$ with probability $1/2$ independently. Let $A$ be the identity nonconformity measure,
$A(z_{1},\dots,z_{m}):=(z_{1},\dots,z_{m}),$
and consider the conformal e-predictor based on $A$ (via (3)). Then the e/variables that it outputs satisfy
$E_{n+1}=\begin{cases}0&\text{with probability $1/2$}\\ \frac{n+1}{k+1}&\text{with probability $1/2$}\end{cases}$ (5)
given $Z_{1},\dots,Z_{n}$, where $k:=\sum_{i=1}^{n}Z_{i}$. This assumes $(Z_{1},\dots,Z_{n})\neq 0$; if $(Z_{1},\dots,Z_{n})=(0,\dots,0)$, the 0 in (5) should be replaced by 1. Since $E_{1},\dots,E_{n}$ uniquely determine $Z_{1},\dots,Z_{n}$ unless $(E_{1},\dots,E_{n})=(1,\dots,1)$, we still have (5) given $E_{1},\dots,E_{n}$ such that $(E_{1},\dots,E_{n})\neq(1,\dots,1)$. Therefore,
$\mathbb{E}(E_{n+1}\mid E_{1},\dots,E_{n})=\frac{n+1}{2(k+1)}$
provided $(E_{1},\dots,E_{n})\neq(1,\dots,1)$, and the expression on the right-hand side may well be different from 1 (albeit will typically be close to 1). This shows that the e-values are not independent: their conditional expectations may be different from their marginal expectations.
Nevertheless, various properties of ergodicity still hold. For example, the following proposition asserts a simple property of ergodicity for the conformal e-values $e_{n}:=f(x_{1},y_{1},\dots,x_{n},y_{n})$, namely their asymptotic online validity.
Proposition 3.
Suppose the observations $(X_{n},Y_{n})$, $n=1,2,\dots$, are IID and e-variables $E_{n}$ are produced by a bounded conformal e-predictor $f$. Then, in the online prediction protocol,
[…]
We can replace the $\leq$” by “$=$” if $f$ is admissible.
Proposition 3 shows that the long-term time average of the e-values for the true labels is bounded above by 1 almost surely. In this sense they are time-wise e-values.
Proof of Proposition 3.
We follow the proof of [24, Lemma 14] (given as Lemma 3.15 in the first edition of [28]) and use the terminology of [21, Chap. 7].
Let $\mathcal{F}_{n}$ be the $\sigma$-algebra generated by the multiset $\lbag Z_{1},\dots,Z_{n-1}\rbag$ and the observations $Z_{n},Z_{n+1},\dots$. (These $\sigma$-algebras form the exchangeable filtration [17, Sect. 5.6].) For each time horizon $N\in\{2,3,\dots\}$, the stochastic sequence $(E_{n}-1,\mathcal{F}_{n})$, $n=N,\dots,1$, is a bounded supermartingale difference. By Hoeffding’s inequality (see, e.g., [28, Sect. A.6.3]), for any $N\in\{2,3,\dots\}$,
$\mathbb{P}\left\{\frac{1}{N}\sum_{n=1}^{N}(E_{n}-1)\geq\epsilon\right\}\leq e^{-2\epsilon^{2}N/C^{2}},$ (7)
where $C$ is an upper bound on the conformal e-predictor $f$. By the Borel–Cantelli lemma [20, Sect. 2.10] (part (a)), the internal inequality in (7) holds only for finitely many $N$ for a fixed $\epsilon>0$. This implies (6).
If $f$ is admissible, we can also apply the same argument to $1-E_{n}$ in place of $E_{n}-1$. ∎
The equation (7) in the proof of Proposition 3 can also be interpreted directly, and its advantage is that it is non-asymptotic. It implies that, for any time horizon $N\geq 2$,
$\mathbb{P}\left\{\frac{1}{N}\sum_{n=1}^{N}E_{n}\geq 1+\epsilon\right\}\leq e^{-2\epsilon^{2}N/C^{2}},$ (8)
which is another expression for the $e_{n}$ being time-wise e-values: their average is approximately bounded above by 1 with high probability (for small $\epsilon$ and large $N$).
Even if we do not assume that the conformal e-predictor is bounded, we can still claim, e.g., that
$\mathbb{P}\left\{\frac{1}{N}\sum_{n=1}^{N}(E_{n}\wedge N^{1/3})\geq 1+\epsilon\right\}\leq e^{-2\epsilon^{2}N^{1/3}}.$ (9)
The last inequality says that $E_{n}\wedge N^{1/3}$ are approximate time-wise e-values (assuming that $\epsilon$ is small and $N\gg\epsilon^{-6}$), and so $E_{n}$ are approximate time-wise e-values if we regard an e-value of $N^{1/3}$ as large enough for the difference between $N^{1/3}$ and larger e-values to be considered unimportant (such as an e-value of 100 on Jeffreys’s scale [9, Appendix B]). (To derive (9) from (8), just set $C:=N^{1/3}$.)
Figure 1: Illustration for Remark 4
Remark 4.
Let us check that we cannot simply drop the requirement that $E$ be bounded in Proposition 3. Suppose that, for each $n$,
$E_{n}:=\begin{cases}n&\text{with probability $\frac{1}{n}$}\\ 0&\text{with probability $\frac{n-1}{n}$}\end{cases}$
independently of $E_{n+1},E_{n+2},\dots$. (For example, $\lvert\mathbf{X}\rvert=1$, $\mathbf{Y}=\mathbb{R}$, each observation is generated from the same continuous probability measure independently, and the nonconformity e-measure is $n$ at the largest $y_{n}$, assuming it is unique, and 00 elsewhere.) Then the distribution of $E_{1},\dots,E_{N}$ is the product distribution
$\prod_{n=1}^{N}\left(\frac{n-1}{n}\delta_{0}+\frac{1}{n}\delta_{n}\right)$ (10)
(where $\delta_{a}$ is the probability measure concentrated on $\{a\}$), since $E_{N},\dots,E_{1}$ are generated independently. This gives us a product distribution for $E_{1},E_{2},\dots$, namely (10) with $N$ replaced by $\infty$. For any $b\in\{1,2,\dots\}$, by Le Cam’s version of Poisson’s theorem [20, Sect. 3.12], the probability of the event that $b$ of the $E_{n}$ for $n\in[2^{k},2^{k+1})$ will satisfy $E_{n}=n$ tends to $e^{-\lambda}\lambda^{b}/b!$ as $k\to\infty$, where $\lambda:=\ln 2$; therefore, this event will happen infinitely often. Therefore, (6) will be violated almost surely; namely, the limsup in it will be $\infty$ almost surely. Figure 1 illustrates this plotting the average conformal e-value $\frac{1}{N}\sum_{n=1}^{N}e_{n}$ vs $N$ for a wide range of $N$, $N\in\{1,\dots,10^{10}\}$ (with $N$ given on a logarithmic scale).
In Appendix A we will check that Proposition 3 can be strengthened to the following statement of the iterated-logarithm type.
Proposition 5.
Suppose the observations $(X_{n},Y_{n})$, $n=1,2,\dots$, are IID. Then, in the online prediction protocol,
[…]
where $C$ is an upper bound on the conformal e-predictor producing $E_{1},E_{2},\dots$.
In conclusion of this section, let me state what I consider to be the main property of validity for e-conformal prediction (it implies the other properties listed here, although by itself it is not particularly intuitive): $\mathbb{E}(E_{n}\mid\mathcal{F}_{n+1})\leq 1$, where $\mathcal{F}$ is the exchangeable filtration and $E_{n}$ is a conformal e/variable. It was introduced in the proof of Proposition 3 but is worth stating separately.
4 Conditional conformal e-predictors
Conformal prediction and conformal e/prediction satisfy the marginal property of validity that we have discussed so far. In this section we will discuss stronger properties of validity, including what we called object-conditional and label-conditional validity in [28, Sect. 4.6]. Suppose, e.g., we have an algorithm for diagnosing Covid. In object-conditional validity, we divide the objects into separate categories (such as men and women) and require validity inside each of the categories, not just on average. In label-conditional validity, we require validity for each possible label, in this case separate validity for people who have Covid and those who do not. More generally, we can talk about observation-conditional validity, where the categories for which we require separate validity are defined in terms of both objects and labels.
The most standard approach to conditional conformal prediction is Mondrian conformal prediction (see, e.g., [28, Sections 4.6.7–4.6.8]), where the observation space $\mathbf{Z}$ is split into a family, often finite, of disjoint categories, and conformal p-values are computed for each category separately. For a relatively narrow group of conformal predictors, object-conditional validity has been established for overlapping categories by Jung et al. [10]. This has been generalized by Gibbs et al. [6] to what may be called “fuzzy categories”: a category becomes a nonnegative function on the object space $\mathbf{X}$, with categories in the sense of subsets of $\mathbf{X}$ corresponding to the indicator functions on $\mathbf{X}$. In this section we will see that the steps of moving from disjoint to overlapping and then to “fuzzy” categories become very simple and natural in conformal e-prediction.
Formally, we have a finite set, called taxonomy, $\mathcal{K}$ of functions $\kappa:\mathbf{Z}\to[0,\infty)$, and we are only interested in nonconformity e-measures $A$ such that, for all $m\in\{1,2,\dots\}$ and all $z_{1},\dots,z_{m}\in\mathbf{Z}$,
$\frac{\sum_{i=1}^{m}\alpha_{i}\kappa(z_{i})}{\sum_{i=1}^{m}\kappa(z_{i})}\leq 1,\quad\forall\kappa\in\mathcal{K},$ (11)
where $(\alpha_{1},\dots,\alpha_{m}):=A(z_{1},\dots,z_{m})$, and $/0$ 0 is interpreted as $1$ when it occurs on the left-hand side. We will say that such an $A$ and the corresponding conformal e-predictor are $\mathcal{K}$-conditional. An important special case, a group-wise taxonomy, is where all $\kappa\in\mathcal{K}$ take values in $\{0,1\}$. In this case we regard $\{z\in\mathbf{Z}:\kappa(z)=1\}$ as the categories. In general, we can still regard $\kappa$ as fuzzy categories.
The requirement (11) formalizes the conditional validity of the conformal e-predictor. But we do not have efficiency inside a category $\kappa$ if (11) holds as “$<$” rather than “$=$”: if we are interested in this category only, we can improve confidence of predictions without violating validity for this category (but perhaps violating validity for other categories). Let us say that a conformal e-predictor is $\mathcal{K}$-exact if it is based on a nonconformity e-measure (also called $\mathcal{K}$-exact) satisfying (11) with $\leq$” replaced by “$=$” for all $m$ and $z_{1},\dots,z_{m}$. Being exact is stronger than what is usually called admissible in decision theory, but exact conformal e-predictors are ideal in the sense of achieving validity without being strictly dominated inside each category.
The allowed nonconformity vectors $(\alpha_{1},\dots,\alpha_{m})$ for a $\mathcal{K}$-exact nonconformity e-measure form the intersection of an affine space and the simplex
$\left\{(\alpha_{1},\dots,\alpha_{m})\in[0,1]^{m}:\alpha_{1}+\dots+\alpha_{m}=m\right\},$
and its dimension is at least $m-1-\lvert\mathcal{K}\rvert$. This affine space can be defined as the shift by the vector $(1,\dots,1)\in\mathbb{R}^{m}$ of the orthogonal complement of the vectors
$\left(\kappa(z_{i})\right)_{i=1}^{m},\quad\kappa\in\mathcal{K}.$ (12)
(Equivalently, we can define the allowed nonconformity vectors $(\alpha_{1},\dots,\alpha_{m})$ as the intersection of the nonnegative orthant in $\mathbb{R}^{m}$ and the orthogonal complement of the vectors (12) extended by adding $(1,\dots,1)\in\mathbb{R}^{m}$ and shifted by $(1,\dots,1)\in\mathbb{R}^{m}$.)
Validity results are easy to state for group-wise taxonomies. First let us state the result in the space domain generalizing Proposition 1.
Proposition 6.
Let $\mathcal{K}$ be a group-wise taxonomy and $f$ be a $\mathcal{K}$-exact conformal e-predictor. For any $n$ and any $\kappa\in\mathcal{K}$, if $Z_{1},\dots,Z_{n},Z$ are IID (or exchangeable),
[…]
Proof.
Let us fix $\kappa\in\mathcal{K}$ and check that (13) is true even conditionally on the $\sigma$-algebra generated by $\kappa(Z_{1}),\dots,\kappa(Z_{n}),\kappa(Z)$, by all $Z_{i}$, $i=1,\dots,n$, with $\kappa(Z_{i})=0$, and by the multiset $B$ consisting of $Z$ and all $Z_{i}$, $i=1,\dots,n$, with $\kappa(Z_{i})=1$. Since, conditionally on $\mathcal{F}$ and inside the event $\kappa(Z)=1$, all orderings of $B$ are equiprobable almost surely, the conditional expectation of $f(Z_{1},\dots,Z_{n},Z)$ is 1 given $\mathcal{F}$ and $\kappa(Z)=1$, which implies (13). ∎
Now we state a validity result in the time domain.
Proposition 7.
Let $\mathcal{K}$ be a group-wise taxonomy. Suppose the observations $Z_{n}$, $n=1,2,\dots$, are IID. Then, in the online prediction protocol,
[…]
for a bounded $\mathcal{K}$-exact conformal e-predictor producing conformal e-variables $E_{n}$.
The event $A\Rightarrow B$ in (14) is defined, as usual, as the union of $B$ and the complement of $A$; therefore, $A\Rightarrow B$ holds almost surely if the event $A\setminus B$ is null.
Proof of Proposition 7.
By the strong law of large numbers, the antecedent of (14) holds with probability 0 or 1, depending on whether $\mathbb{P}(\kappa(Z)=1)=0$ or $\mathbb{P}(\kappa(Z)=1)>0$, where $Z$ is any of $Z_{n}$. In the case $\mathbb{P}(\kappa(Z)=1)=0$, (14) holds vacuously, so let us assume $\mathbb{P}(\kappa(Z)=1)>0$. Our goal is to prove that the consequent of (14) holds almost surely.
By the strong law of large numbers,
[…]
as $N\to\infty$. As in the proof of Proposition 3, we can obtain
[…]
Combining (15) and (16) gives the consequent of (14) holding almost surely. ∎
Remark 8.
The proofs of Propositions 6 and 7 show that their assumptions (exchangeability or being IID) can be weakened, similarly to the case of Mondrian conformal prediction [28, Sect. 11.3.6]. On the other hand, the conclusion of Proposition 7 can be strengthened to give an iterated-logarithm result along the lines of Proposition 5.
5 Split conformal e-predictors
The versions of conformal e-predictors discussed so far are computationally feasible for a large training set only for a narrow class of nonconformity e-measures. The ideas of split conformal e-predictors, discussed in this section, and cross-conformal e-predictors, discussed in the next one, make it possible to extend greatly the practical applicability of conformal e-prediction.
Let us fix a measurable space $\Sigma$ (a summary space). A $\Sigma$-valued split nonconformity measure is a measurable function $A:\mathbf{Z}^{+}\to\Sigma$. Intuitively, $A(z_{1},\dots,z_{m},z)$ encodes how well $z$ conforms to $z_{1},\dots,z_{m}$. A normalizing transformation $N:\Sigma^{+}\to[0,\infty)^{+}$ is an equivariant measurable function that maps every non-empty finite sequence $(\sigma_{1},\dots,\sigma_{m})$ of elements of $\Sigma$ to a finite sequence $(\alpha_{1},\dots,\alpha_{m})$ of the same length of nonnegative numbers whose average is at most 1 (i.e., satisfying (1)). It is admissible if (1) holds with “$=$”.
To apply split conformal e-prediction to a training set $z_{1},\dots,z_{n}$, we split it into two parts, the training set proper $z_{1},\dots,z_{n-c}$ and the calibration set $z_{n-c+1},\dots,z_{n}$. For a new object $x$ and a potential label $y$ for it, we set
$f(z_{1},\dots,z_{n},x,y):=\alpha^{y}$ (17)
where $\alpha^{y}$ is defined using the following steps:
$\sigma_{i}$ $:=A(z_{1},\dots,z_{n-c},z_{n-c+i}),\quad i=1,\dots,c,$
$\sigma^{y}$ $:=A(z_{1},\dots,z_{n-c},(x,y)),$
$(\alpha_{1}^{y},\dots,\alpha_{c}^{y},\alpha^{y})$ $:=N(\sigma_{1},\dots,\sigma_{c},\sigma^{y}).$
For many choices of $A$ and $N$, the split conformal e-predictor (17) will be computationally efficient; this is the case when:
  1. 1.
    Processing the training set proper only once, we can find an easily computable rule transforming $z$ into $A(z_{1},\dots,z_{n-c},z)$.
  2. 2.
    The normalizing transformation $N$ is easily computable.
To give examples of such easily computable $A$ and $N$, suppose we have chosen a suitable learning architecture, such as neural networks, and a way of training it. In the case of pattern recognition, a trained neural network implements a function $F:\mathbf{X}\to\mathbf{P}(\mathbf{Y})$, where $\mathbf{P}(\mathbf{Y})$ is the set of all probability measures on $\mathbf{Y}$, assumed finite and equipped with the discrete $\sigma$-algebra. Given a test object $x\in\mathbf{X}$, this neural network outputs a probability forecast $F(x)\in\mathbf{P}(\mathbf{Y})$ for its label: the true label of $x$ is $y$ with probability $F(x)(\{y\})$. An example of an easily computable (at the prediction stage) rule $A$ is
$A(z_{1},\dots,z_{n-c},(x,y)):=\frac{1}{F_{z_{1},\dots,z_{n-c}}(x)(\{y\})},$ (18)
where $F_{z_{1},\dots,z_{n-c}}:\mathbf{X}\to\mathbf{P}(\mathbf{Y})$ is the neural network found from $z_{1},\dots,z_{n-c}$ as training set. (Training might take a long time, but applying to a new object $x$ is typically quick.) And an example of an easily computable normalizing transformation is
$(\sigma_{1},\dots,\sigma_{m})\mapsto\frac{m}{\sum_{i=1}^{m}\sigma_{i}}(\sigma_{1},\dots,\sigma_{m})$
(cf. (3)), where the summary space is supposed to be $\Sigma\subseteq[0,\infty)$.
Proposition 1, our statement of validity, continues to hold for split conformal e-predictors.
Proposition 9.
For any split conformal e-predictor $f$ and any $n$, if $Z_{1},\dots,Z_{n},(X,Y)$ are exchangeable, we have (4) (with “$=$” if $f$ is admissible).
Proof.
It suffices to notice that (4) holds even conditionally on knowing the observations $Z_{1},\dots,Z_{n-c}$ and the multiset $\lbag Z_{n-c+1},\dots,Z_{n},(X,Y)\rbag$ since all orderings of $\lbag Z_{n-c+1},\dots,Z_{n},(X,Y)\rbag$ are equiprobable almost surely. ∎
6 Cross-conformal e-predictors
Split conformal e-predictors are often computationally efficient, but their predictive efficiency (to be discussed in detail in the next section) may suffer as compared with “full” conformal predictors discussed in Sections 24, since the latter may be said to use the full training set both as training set proper and as calibration set. The idea behind cross-conformal e-prediction is to combine several split conformal predictors in order to achieve better predictive efficiency.
A $\Sigma$-valued split nonconformity measure $A$ is a $\Sigma$-valued cross-nonconformity measure if $A(z_{1},\dots,z_{m},z)$ does not depend on the order of its first $m$ arguments. Given such an $A$ and a normalizing transformation $N$, the corresponding cross-conformal e-predictor is defined as follows. The training sequence $z_{1},\dots,z_{n}$ is randomly split into $K$ non-empty multisets (folds) $z_{S_{k}}$, $k=1,\dots,K$, of equal (or as equal as possible) sizes $\left|S_{k}\right|$, where $K\in\{2,3,\dots\}$ is a parameter of the algorithm, $(S_{1},\dots,S_{K})$ is a partition of the index set $\{1,\dots,n\}$, and $z_{S_{k}}$ consists of all $z_{i}$, $i\in S_{k}$. For each $k\in\{1,\dots,K\}$ and each potential label $y\in\mathbf{Y}$ of the new object $x$, find the output $\alpha_{k}$ of the split conformal e-predictor (based on $A$ and $N$) on the new object $x$ and its postulated label $y$ with $z_{S_{-k}}$ as training set proper and $z_{S_{k}}$ as calibration set, where $S_{-k}:=\cup_{j\neq k}S_{j}=\{1,\dots,n\}\setminus S_{k}$ is the complement to $S_{k}$ (and so $z_{S_{-k}}$ is the complement to the fold $z_{S_{k}}$). The corresponding cross-conformal e-predictor is defined by
$f(z_{1},\dots,z_{n},x,y):=\frac{1}{K}\sum_{k=1}^{K}\alpha_{k}.$
(A slight modification, still provably valid, of this definition is where the arithmetic mean is replaced by the weighted mean with the weights proportional to the sizes $\left|S_{k}\right|$ of the folds.)
Proposition 1 still holds for cross-conformal e-predictors.
Proposition 10.
For any cross-conformal e-predictor $f$ and any $n$, if $Z_{1},\dots,Z_{n},(X,Y)$ are exchangeable, we have (4) (with “$=$” if $f$ is admissible).
Proof.
This follows from Proposition 9 and the arithmetic mean of e-variables being an e-variable (this is obvious and discussed in detail in [32, Sect. 3]). ∎
Remark 11.
To compare informally the outputs of cross-conformal predictors [28, Sect. 4.4] and cross-conformal e-predictors, we can use the rough transformation discussed in [32, Remark 2.3]: a p-value of $p$ roughly corresponds to an e-value of $1/p$. Under this transformation, the arithmetic average of e-values corresponds to the harmonic average of p-values, and the harmonic average is always less than or equal to the arithmetic average [8, Theorem 16]. This suggests that cross-conformal e-prediction produces better results than cross-conformal prediction does. In the opposite direction, the arithmetic average of p-values corresponds to the harmonic average of e-values, which again suggests that cross-conformal e-prediction produces better results.
Remark 12.
It is easy to combine the ideas of this section and Sect. 4 to design conditional cross-conformal e-predictors (or combine Sections 5 and 4), but we stick to the simplest cases.
Remark 13.
Proposition 3 continues to hold for cross-conformal e-predictors and, therefore, it gives its time-wise property of validity in the online mode. However, intuitively, using cross-conformal e-prediction in the online mode defeats the purpose of cross-conformal prediction: once we process $x_{1},y_{1},\dots,x_{n},y_{n}$, we would like to apply the rule that we have found (see item 1 on p. 1) to a large number of new objects rather than just one. There are more complicated settings of “weak teachers” (along the lines of [28, Sect. 3.3]) that combine cross-conformal e-prediction with time-wise validity in useful ways, but we will not discuss them further in this paper.
7 Predictive efficiency of conformal e-predictors
So far we have concentrated on the validity of conformal e-predictors. A valid e/predictor is not allowed to output consistently large e-values for the true labels; namely, the expectation of the e-value for the true label should not exceed 1. On the other hand, for the other labels (we will call them false labels), we would like their e-values to be as large as possible, and this (informal) desideratum is known as efficiency (or predictive efficiency, if there is a risk of confusion with computational efficiency). The topic of this section is ways of measuring the efficiency for conformal e-prediction.
We developed suitable criteria of efficiency for conformal prediction in [28, Sect. 3.1] (whose notation we will use in this section, except that $e$ will stand for e-values). The idea is that a criterion of efficiency cannot be regarded as suitable if in the limiting case of infinite training and test sets it leads to very unnatural optimal nonconformity measures. An example of such an unsuitable criterion of efficiency for conformal prediction is the use of average confidence and credibility, as defined in the first edition of [28] (analogously to our definition in Sect. 2 above) and analysed in [28, Sections 3.1.6–3.1.7]. This section proposes natural criteria of efficiency for conformal e-prediction that lead to natural nonconformity e-measures.
Suppose we have a training set $z_{1},\dots,z_{n}\in\mathbf{Z}$, and we are given a test set $z_{n+1},\dots,z_{n+k}$, where $z_{i}=(x_{i},y_{i})$ for all $i$. Let $e_{i}^{y}:=f(z_{1},\dots,z_{n},x_{i},y)$ be the e-value computed by a given conformal e-predictor $f$ for a postulated label $y$ for a test object $x_{i}$, $i\in\{n+1,\dots,n+k\}$. The average sum
$\frac{1}{k}\sum_{i=n+1}^{n+k}\sum_{y\neq y_{i}}\ln e_{i}^{y}$ (19)
of the log e-values for the false test labels may serve as a measure of the predictive efficiency of $f$ on the test set; we would like it to be as large as possible. Let us call it the observed log criterion of efficiency (the modifier “observed” will be discussed in Remark 15). The expression (19) is natural insofar as averaging logarithms of e-values is ubiquitous in numerous contexts: see, e.g., the five reasons given in [7, (7) and Sect. 3.1].
Let us find the optimal nonconformity e-measure for the observed log criterion in the “idealised setting” akin to the one considered in [28, Sect. 3.1.4] in the case of conformal prediction. For that it will be convenient to modify slightly the definition of a nonconformity e-measure.
An equivalent definition of a nonconformity e-measure $A$ is as a measurable function mapping a nonempty multiset $\lbag z_{1},\dots,z_{m}\rbag$, for any $m\in\{1,2,\dots\}$, and its element $z_{i}$, $i\in\{1,\dots,m\}$, to a nonnegative number satisfying
$\sum_{i=1}^{m}A(\lbag z_{1},\dots,z_{m}\rbag,z_{i})\leq 1$
for all $m\in\{1,2,\dots\}$ and all multisets $\lbag z_{1},\dots,z_{m}\rbag$ of size $m$. The corresponding conformal e-predictor is then defined by
$f(z_{1},\dots,z_{n},x,y):=A(\lbag z_{1},\dots,z_{n},(x,y)\rbag,(x,y)).$ (20)
It is clear that this definition is equivalent to our original definition (2).
An idealised nonconformity e-measure is a function $A:\mathbf{P}(\mathbf{Z})\times\mathbf{Z}\to[0,\infty)$ such that $A(Q,z)$ is measurable in $z\in\mathbf{Z}$ and $\int A(Q,z)Q(\mathrm{d}z)\leq 1$ for any $Q\in\mathbf{P}(\mathbf{Z})$. The intuition behind this definition is that $Q$ represents an infinite training set, which becomes, once the order of its elements is forgotten, the probability distribution of the data. The corresponding idealised conformal e-predictor is $f(Q,x,y):=A(Q,x,y)$ according to (20) (adding another observation to an infinite training set does not change anything); the difference between a nonconformity e-measure and the corresponding conformal e-predictor disappears in the limit.
As in [28, Sect. 3.1.5], we assume that the object space $\mathbf{X}$ and the label space $\mathbf{Y}$ are finite (the latter simply means that we are interested in pattern recognition). We identify a probability measure $Q$ on a finite set $A$ (such as $\mathbf{Y}$) with a function mapping $a\in A$ to $Q(\{a\})$, often omitting the curly braces in expressions such as $Q(\{a\})$. If $Q$ is a probability measure on $\mathbf{Z}=\mathbf{X}\times\mathbf{Y}$, let $Q_{\mathbf{X}}\in\mathbf{P}(\mathbf{X})$ be its marginal probability measure on $\mathbf{X}$, and let $Q_{x}\in\mathbf{P}(\mathbf{Y})$, $x\in\mathbf{X}$, be the conditional probability measure on $\mathbf{Y}$ given $x$:
$Q_{\mathbf{X}}(x):=Q(\{x\}\times\mathbf{Y}),\quad Q_{x}(y):=\frac{Q(x,y)}{Q_{\mathbf{X}}(x)}.$
Let us fix the true data-generating probability measure $P\in\mathbf{P}(\mathbf{Z})$, representing a given infinitely large training set. For simplicity, let us assume that $P(z)>0$ for all $z\in\mathbf{Z}$; this makes $P_{x}(y)$ well-defined and positive for all $x\in\mathbf{X}$ and $y\in\mathbf{Y}$. To formalize the test set being infinitely large as well, we replace the problem of maximizing (19) by the idealised optimization problem
$\int_{\mathbf{Z}}\sum_{y^{\prime}\in\mathbf{Y}\setminus\{y\}}\ln f(P,x,y^{\prime})P(\mathrm{d}(x,y))\to\max,$ (21)
$f=A$ ranging over the conformal e-predictors (equivalently, over the nonconformity e-measures). This corresponds to letting $k\to\infty$ in (19).
Proposition 14.
The optimal nonconformity e-measure $A$ under the observed log criterion (21) is given by the odds
$A(P,x,y):=\frac{1}{\lvert\mathbf{Y}\rvert-1}\frac{1-P_{x}(y)}{P_{x}(y)}$ (22)
against the true label being $y$ conditional on the object $x$.
The rest of this section will be mainly devoted to the proof of Proposition 14. First let us check that (22) is indeed a nonconformity e-measure. It is even even true that its average is 1 over each $P_{x}$:
$\int A(P,x,y)P_{x}(\mathrm{d}y)=\frac{1}{\lvert\mathbf{Y}\rvert-1}\sum_{y\in\mathbf{Y}}(1-P_{x}(y))=1.$
We start from the case where we have no objects, corresponding to $\lvert\mathbf{X}\rvert=1$ (the only object does not carry any information). This case is not only a gentle introduction to the general case, but also corresponds to the situation where our prediction is fully conditional on the object $x$ (remember that earlier we assumed $\lvert\mathbf{X}\rvert<\infty$, and so the full conditioning is feasible for an infinite training set). Now we can regard $P$ to be a probability measure on the label space $\mathbf{Y}$, and our goal is to find the optimal e-variable $E=Q/P$, where $Q$ is an alternative probability measure on $\mathbf{Y}$.
In the case $\lvert\mathbf{X}\rvert=1$, the optimization problem (21) is
$\int\ln\frac{Q}{P}\,\mathrm{d}1-\int\ln\frac{Q}{P}\,\mathrm{d}P\to\max,$
where $1$ is the counting measure on $\mathbf{Y}$ ($11(y):=$ 1 for all $y\in\mathbf{Y}$). This optimization problem is equivalent to
$\int\ln Q\,\mathrm{d}(1-P)\to\max,$
which gives $Q(y)\propto 1-P(y)$. So the optimal e-values are
$E(y)\propto\frac{1-P(y)}{P(y)};$
i.e., $E(y)$ is proportional to the odds against observing label $y$. The full expression is
$E(y)=\frac{1}{\lvert\mathbf{Y}\rvert-1}\frac{1-P(y)}{P(y)},$
which agrees with (22).
Now let us get rid of the assumption $\lvert\mathbf{X}\rvert=1$. We will apply the result of the previous paragraph to each $P_{x}$, $x\in\mathbf{X}$. To find an explicit expression for the optimal e-variable $E:\mathbf{Z}\to[0,\infty)$, we split it into $E_{x}:\mathbf{Y}\to[0,\infty)$, $x\in\mathbf{X}$, defined by $E_{x}(y):=E(x,y)$. Our optimization problem is
$\int_{\mathbf{X}}\int_{\mathbf{Y}}\ln E_{x}\,\mathrm{d}(1-P_{x})P_{\mathbf{X}}(\mathrm{d}x)\to\max$ (23)
subject to the constraint
$\int_{\mathbf{X}}\int_{\mathbf{Y}}E_{x}\,\mathrm{d}P_{x}P_{\mathbf{X}}(\mathrm{d}x)\leq 1;$ (24)
without loss of generality, we replace $\leq$” by “$=$” in (24). Now the constraint (24) can be rewritten as
$\int_{\mathbf{Y}}E_{x}\,\mathrm{d}P_{x}=1+\gamma_{x},\quad\int_{\mathbf{X}}\gamma_{x}P_{\mathbf{X}}(\mathrm{d}x)=0,$
where the new variables $\gamma_{x}$ take values in $[-1,\infty)$. By the result of the previous paragraph,
$E_{x}=(1+\gamma_{x})\tilde{E}_{x},$
where
$\tilde{E}_{x}(y)=\frac{1}{\lvert\mathbf{Y}\rvert-1}\frac{1-P_{x}(y)}{P_{x}(y)}$
is the normalized version of $E_{x}$. Maximizing the overall objective function in (23) can be rewritten as
$\int_{\mathbf{X}}\int_{\mathbf{Y}}\ln\tilde{E}_{x}\,\mathrm{d}(1-P_{x})P_{\mathbf{X}}(\mathrm{d}x)+(\lvert\mathbf{Y}\rvert-1)\int_{\mathbf{X}}\ln(1+\gamma_{x})P_{\mathbf{X}}(\mathrm{d}x)\to\max,$
and so our optimization problem reduces to
$\int_{\mathbf{X}}\ln(1+\gamma_{x})P_{\mathbf{X}}(\mathrm{d}x)\to\max$ (25)
under the constraint
$\int_{\mathbf{X}}\gamma_{x}P_{\mathbf{X}}(\mathrm{d}x)=0.$
By Jensen’s inequality, the max in (25) is 0, and it is attained for $\gamma_{x}=0$. This completes the proof of Proposition 14.
Remark 15.
In the terminology of [28, Sect. 3.1], the criterion of efficiency (19) is “observed” in that the efficiency of the conformal e-predictor on a test observation $z_{i}$ is measured by an expression, namely $\sum_{y\neq y_{i}}\ln e_{i}^{y}$, that depends on the observed true label $y_{i}$. A natural alternative to (19) is the prior log criterion
$\frac{1}{k}\sum_{i=n+1}^{n+k}\sum_{y}\ln e_{i}^{y}$ (26)
obtained by replacing $\sum_{y\neq y_{i}}$ by $\sum_{y}$. The criterion (26) is called “prior” since for it the dependence on the true label disappears; we can compute the sum $\sum_{y}$ prior to observing the true label. Its idealised version is
$\int_{\mathbf{X}}\sum_{y^{\prime}\in\mathbf{Y}}\ln f(P,x,y^{\prime})P_{\mathbf{X}}(\mathrm{d}x)\to\max.$
Adapting the argument given above to this idealised criterion (and slightly simplifying the argument), we can see that the optimal nonconformity e-measure under this criterion is
$A(P,x,y):=\frac{1}{\lvert\mathbf{Y}\rvert P_{x}(y)}.$
(We have already used this nonconformity e-measure with $P_{x}$ replaced by its estimate: see (18) above.) Notice that while in conformal prediction suitable (“conditionally proper”) observed and prior criteria of efficiency lead to the same optimal conformity measures [28, Theorem 3.1], in the case of conformal e-prediction the optimal nonconformity e-measures are different, albeit typically very close for a large label space $\mathbf{Y}$.
8 Conformal e-testing
In this and next sections we discuss testing the assumption of exchangeability. We start in this section from batch testing, which is standard in statistics (although we use terminology that is standard in conformal testing rather than statistics). We assume that the number $N$ of observations is fixed and known in advance. At the end of step $N$ we are required to make a decision whether to reject the hypothesis of exchangeability, or at least to present the amount of evidence that we found against the null hypothesis of exchangeability. In the next section we will consider testing the assumption of exchangeability repeatedly, which we will refer to as multistage testing; as soon as we suspect that exchangeability is violated, we raise an alarm, and we are allowed do so more than once.
Conformal e-testing in the batch mode is defined in two steps. First, a function $E:\mathbf{Z}^{N}\to[0,\infty)$ is called a basic conformal e-test if it has the form
$E(z_{1},\dots,z_{N})=\prod_{n=1}^{N}f(z_{1},\dots,z_{n}),\quad\forall(z_{1},\dots,z_{N})\in\mathbf{Z}^{N},$
for some conformal e-predictor $f$. This definition corresponds to the following gambling picture: we start from an initial capital of $1$ and then compound the conformal e-values as usual by multiplying them (ignoring the fact that the $n$th e-value output by $f$ is not an e-value conditionally on $e_{1},\dots,e_{n-1}$). And second, a conformal e-test is defined as a convex combination of basic conformal e-tests: namely, the conformal e-tests are defined as convex mixtures $\lambda_{1}E_{1}+\dots+\lambda_{k}E_{k}$ of basic conformal e-tests $E_{1},\dots,E_{k}$, where $k\in\{1,2,\dots\}$, $\lambda_{i}\in[0,1]$, and $\lambda_{1}+\dots+\lambda_{k}=1$.
Proposition 16.
Suppose $Z_{1},\dots,Z_{N}$ are exchangeable. For any conformal e-test $E$, $E(Z_{1},\dots,Z_{N})$ is a bona fide e-variable. It satisfies $\mathbb{E}(E(Z_{1},\dots,Z_{N}))=1$ if the underlying conformal e-predictor is admissible.
Proof.
Assume, without loss of generality, that $E$ is a basic conformal e-test. Let us set $E_{n}:=f(Z_{1},\dots,Z_{n})$, where $f$ is the underlying conformal e-predictor, and use the exchangeable filtration $(\mathcal{F}_{n})$. We can show that $\mathbb{E}(E_{1}\dots E_{N}\mid\mathcal{F}_{N+1})\leq 1$ a.s. by induction in $N$ (for $N=0$ this statement is vacuously true):
$\mathbb{E}(E_{1}\dots E_{N}\mid\mathcal{F}_{N+1})=\mathbb{E}(\mathbb{E}(E_{1}\dots E_{N}\mid\mathcal{F}_{N})\mid\mathcal{F}_{N+1})\\ =\mathbb{E}(E_{N}\mathbb{E}(E_{1}\dots E_{N-1}\mid\mathcal{F}_{N})\mid\mathcal{F}_{N+1})\leq\mathbb{E}(E_{N}\mid\mathcal{F}_{N+1})\leq 1\quad\text{a.s.}$
If the underlying conformal e-predictor is admissible, both $\leq$” become “$=$”. ∎
Proposition 16 is a statement of validity of conformal e-testing. If $E$ is chosen in advance and takes a very large value on the realized data sequence, we are justified in rejecting the assumption of exchangeability.
As the next step, we explore the efficiency of conformal e-testing along the lines of the treatment of the efficiency of conformal testing in [28, Sect. 9.1]. As in the case of conformal testing [28, Sect. 9.1], in our (rather weak) statement of efficiency we will restrict ourselves to the binary case, $\mathbf{Z}=\{0,1\}$.
First we define unrestricted testing of exchangeability. A probability measure $P$ on $\mathbf{Z}^{N}$ is exchangeable if, for any permutation $\pi$ of $\{1,\dots,N\}$ and any measurable $A\subseteq\mathbf{Z}^{N}$,
$P(A)=P\{(z_{1},\dots,z_{N}):(z_{\pi(1)},\dots,z_{\pi(N)})\in A\}.$
The upper exchangeability probability $\operatorname{\mathbb{P}^{\mathrm{exch}}}$ of measurable $A\subseteq\mathbf{Z}^{N}$ is defined as
$\operatorname{\mathbb{P}^{\mathrm{exch}}}(A):=\sup_{P}P(A),$
$P$ ranging over the exchangeable probability measures on $\mathbf{Z}^{N}$. The intuition behind $\operatorname{\mathbb{P}^{\mathrm{exch}}}(A)$ is that, if it is very small, $A$ can be used for testing the exchangeability of the data-generating distribution: we are entitled to reject exchangeability if $(z_{1},\dots,z_{N})\in A$ provided $A$ is chosen in advance.
The upper conformal e-probability $\operatorname{\mathbb{P}^{\mathrm{ce}}}$ of $A\subseteq\mathbf{Z}^{N}$ is defined as
$\operatorname{\mathbb{P}^{\mathrm{ce}}}(A):=\inf\left\{\alpha:\exists E\;\forall(z_{1},\dots,z_{N})\in A:E(z_{1},\dots,z_{N})\geq 1/\alpha\right\},$
$E$ ranging over the conformal e-tests. The intuition is the same: if $\operatorname{\mathbb{P}^{\mathrm{ce}}}(A)$ is very small, we are entitled to reject exchangeability if we observe $(z_{1},\dots,z_{N})\in A$, again assuming $A$ is chosen in advance. The difference from $\operatorname{\mathbb{P}^{\mathrm{exch}}}(A)$ is that now the lack of exchangeability must be demonstrated via conformal e-testing.
The following proposition is an analogue of [28, Proposition 9.5]. Intuitively, its second statement says that conformal e-testing is universal in the batch mode: if lack of exchangeability can be demonstrated at all, it can be demonstrated (albeit less convincingly) using conformal e-testing.
Proposition 17.
For any event $A\subseteq\mathbf{Z}^{N}$, $\operatorname{\mathbb{P}^{\mathrm{exch}}}(A)\leq\operatorname{\mathbb{P}^{\mathrm{ce}}}(A)$. Assuming $\mathbf{Z}=\{0,1\}$, $\operatorname{\mathbb{P}^{\mathrm{ce}}}(A)\leq N\operatorname{\mathbb{P}^{\mathrm{exch}}}(A)$.
Proof.
The inequality $\operatorname{\mathbb{P}^{\mathrm{exch}}}(A)\leq\operatorname{\mathbb{P}^{\mathrm{ce}}}(A)$ follows immediately from Markov’s inequality applied to conformal e-tests $E$ in combination with Proposition 16: if $E(z_{1},\dots,z_{N})\geq 1/\alpha$ for all $(z_{1},\dots,z_{N})\in A$,
$\operatorname{\mathbb{P}^{\mathrm{exch}}}(A)\leq\mathbb{P}(E(Z_{1},\dots,Z_{N})\geq 1/\alpha)\leq\frac{\mathbb{E}(E(Z_{1},\dots,Z_{N}))}{1/\alpha}\leq\alpha,$
assuming $Z_{1},\dots,Z_{N}$ are exchangeable.
Now assume $\mathbf{Z}=\{0,1\}$; this part of the proof will be a modification of the proof of Proposition 9.5 in [28, Sect. 9.4.2]. As a first step, notice that it suffices to prove $\operatorname{\mathbb{P}^{\mathrm{ce}}}(A)\leq\operatorname{\mathbb{P}^{\mathrm{exch}}}(A)$ for any nonempty $A\subseteq\mathbf{Z}^{N}$ such that each sequence in $A$ has the same number of 1s. Let us fix such an $A$, and let $K\in\{0,\dots,N\}$ be the number of 1s in the elements of $A$. For each sequence $\zeta=(z_{1},\dots,z_{N})\in A$, consider the basic conformal e-test $E_{\zeta}=e_{1}\dots e_{N}$, where the $n$th e-value $e_{n}$ is
$e_{n}:=\begin{cases}n/k&\text{if $z_{n}=1$}\\ n/(n-k)&\text{if $z_{n}=0$},\end{cases}$
$k$ being the number of 1s among the first $n$ elements of $\zeta$. (This was called “reckless gambling” in [28, end of Sect. 9.1.2].) The product $E_{\zeta}=e_{1}\dots e_{N}$ will then be
$\frac{N!}{K!(N-K)!}=\binom{N}{K}.$
The arithmetic mean of $E_{\zeta}$ over $\zeta\in A$ witnesses that
$\operatorname{\mathbb{P}^{\mathrm{ce}}}(A)\leq\lvert A\rvert/\binom{N}{K}=\operatorname{\mathbb{P}^{\mathrm{exch}}}(A).\qed$
9 Multistage exchangeability testing
In the previous section we discussed testing the assumption of exchangeability once, but in some important applications we would like to test it repeatedly over time in an online mode (see, e.g., [28, Sect. 8.3]). As usual, such procedures are required to satisfy properties of validity and efficiency. Validity means that, under exchangeability, the probability or frequency of alarms (which in this case are false alarms) should be bounded by a prespecified constant. Efficiency means that, if exchangeability is violated at some point, an alarm should be raised as quickly as possible afterwards.
In this paper we concentrate on validity of multistage testing; as typical for conformal prediction, we do not have enough structure for efficiency results, unless they are empirical or asymptotic or very weak (as those in the previous section). See [28, Sect. 8.6.3] for a more detailed discussion.
Suppose we observe a sequence of e-values $e_{1},e_{2},\dots$ output by a conformal e-predictor in the online protocol, as described in Sect. 3, under exchangeability. The CUSUM procedure [13] raises the $k$th alarm, $k=1,2,\dots$, at the time
$\tau_{k}:=\min\left\{n>\tau_{k-1}:\max_{i\in\{\tau_{k-1}+1,\dots,n\}}e_{i}\dots e_{n}\geq c\right\},$ (27)
where $\tau_{0}:=0$ and $c>1$ is the parameter of the procedure. It is usually applied in the situation where the observations $z_{i}$ are generated independently first from a known probability measure $Q_{0}\in\mathbf{P}(\mathbf{Z})$ and then from another known probability measure $Q_{1}\in\mathbf{P}(\mathbf{Z})$, and where $e_{i}$ is the likelihood ratio of $Q_{1}$ to $Q_{0}$ evaluated at $z_{i}$ (see, e.g., [16, Sect. 6.2]). In our current context, where the $e_{i}$ are conformal e-values, we may call it the conformal CUSUM e/procedure.
The following proposition gives an asymptotic property of validity for the conformal CUSUM e-procedure.
Proposition 18.
Let $A_{n}$ be the number of alarms
$A_{n}:=\max\{k\mid\tau_{k}\leq n\}$ (28)
raised by the conformal CUSUM e-procedure (27) after processing the first $n$ observations $Z_{1},\dots,Z_{n}$. Then
[…]
provided the observations $Z_{1},Z_{2},\dots$ are exchangeable.
A popular modification of the CUSUM procedure is the Shiryaev–Roberts procedure, which replaces the $\max$ in (27) by the sum $\sum$ (see, e.g., [16, Sect. 6.5]). To state our next validity result (Proposition 19 below), we need the modification
$\tau_{k}:=\min\left\{n>\tau_{k-1}:\max_{i\in\{\tau_{k-1}+1,\dots,n\}}\sum_{j=i}^{n}e_{i}\dots e_{j}\geq c\right\},$ (30)
which we will call the reverse Shiryaev–Roberts procedure (or conformal reverse Shiryaev–Roberts e-procedure to emphasize $e_{i}$ being produced by a conformal e-predictor).
Proposition 19.
Let $A_{n}$ be defined by (28) for the conformal reverse Shiryaev–Roberts e-procedure (30). Then we still have (29) provided the observations $Z_{1},Z_{2},\dots$ are exchangeable.
See Appendix B for the proofs. We will, of course, deduce Proposition 18 from Proposition 19: it is clear that the Shiryaev–Roberts procedure (regular or reverse) raises alarms at least as often as CUSUM does. However, the following example shows an advantage of CUSUM from an intuitive point of view.
Example 20.
Consider the vacuous admissible conformal e-predictor identically equal to 1. The CUSUM procedure based on it will never raise alarms, while the Shiryaev–Roberts e-procedure will raise alarms every $\lceil c\rceil$th step, thereby fully using the permission in the target property of validity (29). This example shows one feature of the Shiryaev–Roberts procedure (shared by the reverse Shiryaev–Roberts procedure) that can be considered its disadvantage: while the CUSUM procedure raises an alarm when it has genuine evidence for disorder, the Shiryaev–Roberts procedure may raise an alarm simply because it is allowed to do so by a given constraint; we might not have any evidence for disorder.
10 Conclusion
In this paper we have discussed (even if some of them very briefly) four strengths of conformal prediction:
  1. 1.
    As set predictors, conformal predictors possess a very simple property of validity: their probability of error is bounded by a prespecified constant. This advantage is lost in conformal e-prediction (although conformal e-predictors enjoy other properties of validity).
  2. 2.
    In the online prediction protocol, conformal predictors as set predictors make errors at different steps independently. (Conformal e-predictors satisfy a weakened version of this property, as discussed in Sect. 3.)
  3. 3.
    In the online prediction protocol, conformal prediction produces bona fide test martingales, which can then be used in one-off and multistage conformal testing. (Weakened variants for conformal e-prediction are discussed in Sections 8 and 9.)
  4. 4.
    In the case of regression, conformal prediction can be used for producing conformal predictive distributions, and so conformal predictors can also be used as probabilistic predictors. In this role, conformal predictors are automatically well-calibrated. This advantage is lost for conformal e-prediction.
Two of these strengths, 1 and 4, appear to be clear advantages of conformal prediction over conformal e-prediction. For strengths 2 and 3, this is less obvious, but the picture for conformal prediction still appears simpler and nicer.
Conformal e-prediction has at least two advantages of its own:
  • Designing conditional conformal e-predictors is much easier and using e-values adds flexibility, especially as compared with Mondrian conformal predictors; see Sect. 4.
  • Cross-conformal e-predictors are provably valid, unlike cross-conformal predictors, as discussed in Sect. 6.
Looking for further advantages is an interesting direction of further research.
Acknowledgments
Many thanks to Aaditya Ramdas for pointing out serious deficiencies in presentation in an earlier version. This research has been partially supported by Astra Zeneca, Stena Line, and Mitie.
Appendix A Law of the iterated logarithm for e-flows
The main goal of this appendix is to prove the law of the iterated logarithm in the form of Proposition 5. As I could not find this version of the law of the iterated logarithm in literature, I will state it in a general form in this appendix.
Let $\mathcal{F}_{n}$, where $n\in\mathbb{Z}$ and $\mathbb{Z}$ stands for the set of integer numbers, be an increasing sequence of $\sigma$-algebras on a given sample space $\Omega$,
$\dots\subseteq\mathcal{G}_{-1}\subseteq\mathcal{G}_{0}\subseteq\mathcal{G}_{1}\subseteq\dots,$
and $\xi_{n}$, $n\in\mathbb{Z}$, be an adapted sequence of random variables, meaning that each $\xi_{n}$ is $\mathcal{G}_{n}$-measurable. Such an adapted sequence is an e-flow if it is nonnegative and
$\mathbb{E}(\xi_{n}\mid\mathcal{G}_{n-1})\leq 1,\quad n\in\mathbb{Z}.$ (31)
Let us say that the e-flow is exact if (31) holds with “$=$” in place of $\leq$”.
The main example of an e-flow in the context of this paper is
$\mathcal{G}_{n}:=\begin{cases}\mathcal{F}_{-n}&\text{if $n<0$}\\ \mathcal{F}_{1}&\text{otherwise},\end{cases}$
where $\mathcal{F}_{n}$, $n=1,2,\dots$, is the exchangeable filtration (introduced in Sect. 3: $\mathcal{F}_{n}$ is generated by $\lbag Z_{1},\dots,Z_{n-1}\rbag,Z_{n},Z_{n+1},\dots$), and
$\xi_{n}:=\begin{cases}E_{-n}&\text{if $n<0$}\\ 1&\text{otherwise},\end{cases}$
where $E_{n}$ is the $n$th conformal e-variable, $E_{n}:=f(Z_{1},\dots,Z_{n})$. Let me call this the conformal e-flow. The fact that the adapted sequence $(\xi_{n},\mathcal{G}_{n})$ defined in this way is an e-flow is the main property of validity of conformal e-prediction. The main part of the conformal e-flow is $(\xi_{n},\mathcal{G}_{n})$ for $n<0$; the extension to all integer $n$ is vacuous. Exact e-flows correspond to admissible conformal predictors.
There are two natural laws of the iterated logarithm for e-flows: the forward law describes the behaviour, as $N\to\infty$, of the sums $\sum_{n=1}^{N}(\xi_{n}-1)$, and the backward law describes the behaviour of the sums $\sum_{n=-N}^{-1}(\xi_{n}-1)$ The forward law is just a restatement of the standard martingale law of the iterated logarithm, since with each forward sequence $\xi_{n}$, $n\geq 1$, we can associate a supermartingale that carries the same information by setting
$X_{n}:=\sum_{i=1}^{n}(\xi_{i}-1),$ (32)
with $X_{0}:=0$; remember that the defining property of being a supermartingale is $\mathbb{E}(X_{n}\mid\mathcal{F}_{n-1})=X_{n-1}$. Applying the standard law of the iterated logarithm to the supermartingale $X_{n}$ gives us the following corollary.
Corollary 21.
For any bounded e-flow $(\xi_{n},\mathcal{G}_{n})$,
[…]
where $C>1$ is an upper bound for $\xi_{n}$, and there exists an e-flow $(\xi_{n},\mathcal{G}_{n})$ bounded above by $C$ (assuming $C>1$) such that
[…]
(Alternatively, the proof of the backward law of the iterated logarithm given later in this appendix will also prove the forward law, after trivial modifications.) The assumption $C>1$ only excludes a trivial case.
Such a reduction to the supermartingale case is impossible for the backward law. Luckily, however, the standard proof of the law of the iterated logarithm can be easily adapted to the backward case and allows us to establish the following backward law of the iterated logarithm.
Proposition 22.
For any bounded e-flow $(\xi_{n},\mathcal{G}_{n})$, we have
[…]
where $C>1$ is an upper bound, and there exists an e-flow $(\xi_{n},\mathcal{G}_{n})$ bounded above by $C$ (assuming $C>1$) that satisfies
[…]
The proof of Proposition 22 uses the standard basic scheme (see, e.g., [21, Sect. 4.4]). As a first step we fix an e-flow $(\xi_{n},\mathcal{G}_{n})$ and assume, without loss of generality, that it is exact (an easy way to see that there is no loss of generality is to apply the idea of coupling [4, Sect. 7.4]).
We need an auxiliary exponential supermartingale from Stout’s proof of the law of the iterated logarithm for martingales [22, Lemma 5.4.1].
Lemma 23.
Let $\eta_{1},\eta_{2},\dots$ be a martingale difference bounded above by a constant $c$, $\eta_{n}\leq c$, w.r. to a filtration $\mathcal{F}_{0},\mathcal{F}_{1},\dots$. Set $S_{n}:=\sum_{i=1}^{n}\eta_{i}$. Let $\delta\in(0,1/c]$ be another constant. Then
$T_{n}:=\exp\left(\delta S_{n}-\frac{\delta^{2}}{2}\left(1+\frac{\delta c}{2}\right)\sum_{i=1}^{n}\mathbb{E}(\eta_{i}^{2}\mid\mathcal{F}_{i-1})\right)$ (35)
is a supermartingale.
In fact, Stout’s lemma only assumes that $\eta_{n}$ is a supermartingale difference, but we do not need this generality. Let us derive a corollary of this lemma that will allow us to establish (33). First we notice that $\mathbb{E}(\eta_{i}^{2}\mid\mathcal{F}_{i-1})\leq c$ (the largest $\mathbb{E}(\eta_{i}^{2}\mid\mathcal{F}_{i-1})=c$ is attained for $\eta_{i}\in\{-1,c\}$ taking value $c$ with probability $1/(c+1)$); this is spelled out in the following lemma.
Lemma 24.
For any $c>0$, $\max_{\eta}\mathbb{E}(\eta^{2})=c$, $\eta$ ranging over the random variables with $\mathbb{E}(\eta)=0$ and $\eta\in[-1,c]$.
Proof.
This is a special case of the Bhatia–Davis inequality [3]: see Theorem 1 there and the remark after its first proof. ∎
Since $T_{n}$ defined by (35) is a test supermartingale, we have, by Ville’s inequality,
$\mathbb{P}\left\{\max_{n=0,\dots,N}T_{n}\geq\gamma\right\}\leq\frac{1}{\gamma}$
for any $\gamma>1$ and any upper bound $N$ on $n$. By Lemma 24 (modified to cover $\mathbb{E}(\eta^{2}\mid\mathcal{F})$ in place of $\mathbb{E}(\eta^{2})$), this implies
$\mathbb{P}\left\{\max_{n=0,\dots,N}S_{n}\geq\frac{\delta}{2}\left(1+\frac{\delta c}{2}\right)cN+\frac{\ln\gamma}{\delta}\right\}\leq\frac{1}{\gamma}.$ (36)
The minimum over $\delta$ of the sum
$\frac{\delta}{2}cN+\frac{\ln\gamma}{\delta}$
(with the term $\frac{\delta c}{2}$ in (36) ignored for now) is attained at
$\delta:=\sqrt{2\ln\gamma/(cN)},$ (37)
and substituting this expression for $\delta$ into (36) gives
$\mathbb{P}\left\{\max_{n=0,\dots,N}S_{n}\geq\sqrt{2cN\ln\gamma}+\frac{c\ln\gamma}{2}\right\}\leq\frac{1}{\gamma}$ (38)
(the condition $\delta\leq 1/c$ will be satisfied when we apply this inequality later in the proof). The inequality (38) is also applicable to $-S_{n}$ in place of $S_{n}$ provided $c\geq 1$, and for any $c>0$ it becomes applicable to $-S_{n}$ in place of $S_{n}$ if we replace the second entry of $c$ in it by 1 and assume $\delta\leq 1$.
Proof of Proposition 22.
Let $\lambda>1$ (later we will also let $\lambda\to 1$) and set $n_{k}:=\lceil\lambda^{k}\rceil$, $k=1,2,\dots$. For a given $k$, we will use the notation, for $k,n\in\{1,2,\dots\}$,
$\eta^{k}_{n}:=\xi_{-n_{k}-1+n}-1\text{ and }S^{k}_{n}:=\sum_{i=1}^{n}\eta_{i},$
so that $\eta^{k}_{n}$, $n=1,2,\dots$, is a martingale difference taking values in $[-1,C-1]$ and $S^{k}_{n}$, $n=1,2,\dots$, is a martingale. Set $c:=C-1$.
Later we will choose a suitable function $\psi:\{1,2,\dots\}\to\mathbb{R}$; roughly, $\psi(n)$ will be an upper bound for $\sum_{i=-n}^{-1}(\xi_{i}-1)$. Let $A_{k}$ be the event that $S^{k}_{n_{k}}\geq\psi(n_{k})$ and $B_{k}$ be the event that $-S^{k}_{n}\geq\psi(n_{k}-n_{k-1})$ for some $n\in(0,n_{k}-n_{k-1}]$. Namely, we will choose $\psi$ in such a way that, for sufficiently large $k$,
$\mathbb{P}(A_{k})\leq\mathbb{P}(S^{k}_{n}\geq\psi(n_{k})\text{ for some $n\leq n_{k}$})\leq k^{-\lambda}$ (39)
and
$\mathbb{P}(B_{k})=\mathbb{P}(-S^{k}_{n}\geq\psi(n_{k}-n_{k-1})\text{ for some $n\leq n_{k}-n_{k-1}$})\leq k^{-\lambda}.$ (40)
By the Borel–Cantelli lemma, as $\sum_{k}\mathbb{P}(A_{k})<\infty$ and $\sum_{k}\mathbb{P}(B_{k})<\infty$, $A_{k}$ and $B_{k}$ will hold only for finitely many $k$.
In order for the inequalities “${}\leq k^{-\lambda}$” in (39) and (40) to hold, we can set, according to (38) (with $-S_{n}$ in place of $S_{n}$ in the case of (40)),
$\psi(n_{k})$ $=\sqrt{2cn_{k}\ln(k^{\lambda})}+\frac{c\ln(k^{\lambda})}{2}\sim\sqrt{2cn_{k}\lambda\ln k}$
$\psi(n_{k}-n_{k-1})$ $=\sqrt{2c(n_{k}-n_{k-1})\ln(k^{\lambda})}+\frac{\ln(k^{\lambda})}{2}\sim\sqrt{2c(n_{k}-n_{k-1})\lambda\ln k}.$
Therefore, we can choose
$\psi(n)\sim\sqrt{2c\lambda n\ln\ln n}.$
The conditions $\delta\leq 1/c$ and $\delta\leq 1$ mentioned earlier indeed hold, from some $k$ on, for (37), $\gamma:=k^{\lambda}$, and $N:=n_{k}$.
According to (39) and (40), we will have, from some $k$ on,
$S^{k}_{n_{k}}$ $\leq\psi(n_{k}),$
$-S^{k}_{n}$ $\leq\psi(n_{k}-n_{k-1})\text{ for all $n\leq n_{k}-n_{k-1}$}.$
For any sufficiently large $N$, these inequalities imply the inequality in
$\sum_{n=-N}^{-1}(\xi_{n}-1)=\sum_{n=1}^{n_{k}}\eta_{n}^{k}-\sum_{n=1}^{n_{k}-N}\eta_{n}^{k}=S^{k}_{n_{k}}-S^{k}_{n_{k}-N}\leq\psi(n_{k})+\psi(n_{k}-n_{k-1})\\ \sim\sqrt{2c\lambda n_{k}\ln\ln n_{k}}+\sqrt{2c\lambda(n_{k}-n_{k-1})\ln\ln(n_{k}-n_{k-1})},$
where $k$ is the value satisfying $N\in[n_{k-1},n_{k})$. Since we can take $\lambda$ arbitrarily close to 1, this completes the proof of (33).
To prove (34) for some bounded e-flow, suppose
$\xi_{n}=\begin{cases}C&\text{with probability $1/C$}\\ 0&\text{with probability $1-1/C$}\end{cases}$
independently for $n=-1,-2,\dots$. Then $\operatorname{var}(\xi_{n})=C-1$, and applying the standard law of the iterated logarithm gives (34). ∎
Our argument given above proves the following more standard law of the iterated algorithm (which we do not need in the paper).
Proposition 25.
Let $\eta_{n}$, $n\in\mathbb{Z}$ be a bounded two-sided sequence of random variables adapted to a filtration $\mathcal{G}_{n}$, $n\in\mathbb{Z}$. Suppose $\mathbb{E}(\eta_{n}\mid\mathcal{G}_{n-1})=0$ for all $n\in\mathbb{Z}$. Set
$S_{N}:=\sum_{n=-N}^{-1}\eta_{n}\text{ and }A_{N}:=\sum_{n=-N}^{-1}\mathbb{E}(\eta^{2}_{n}\mid\mathcal{G}_{n-1}).$ (41)
Then
[…]
provided $A_{N}\to\infty$ a.s. as $N\to\infty$.
In the proof of Proposition 25 we should use the stopping times $\tau_{k}:=\min\{n:A_{n}\geq\lambda^{k}\}$ instead of the constant stopping times $n_{k}$. Of course, this proposition will remain true if we replace (41) by
$S_{N}:=\sum_{n=1}^{N}\eta_{n}\text{ and }A_{N}:=\sum_{n=1}^{N}\mathbb{E}(\eta^{2}_{n}\mid\mathcal{G}_{n-1}),$
but then it will become just a special case of the standard martingale law of the iterated logarithm.
Appendix B Proofs for Sect. 9
The idea of the proof of Proposition 19 consists, as usual, in reversing the direction of time. Let $N$ be a sufficiently large natural number, and let $\mathcal{F}_{n}$, $n\in\{1,\dots,N+1\}$, be the $\sigma$-algebra generated by the multiset $\lbag Z_{1},\dots,Z_{n-1}\rbag$ and observations $Z_{n},\dots,Z_{N}$. We also allow $n=N+1$, in which case $\mathcal{F}_{N+1}$ is the $\sigma$-algebra generated by the multiset $\lbag Z_{1},\dots,Z_{N}\rbag$. This is a finite-horizon version of the exchangeable filtration.
Let $E_{n}$ be the $n$th conformal e-variable (having $e_{n}$ as it value). Then $(E_{n},\mathcal{F}_{n})$, $n=N,\dots,1$, is an exact e-flow over the finite time interval $N,\dots,1$, in the sense
$\mathbb{E}(E_{n}\mid\mathcal{F}_{n+1})=1,\quad n=N,\dots,1.$
The corresponding martingale is $(T_{n},\mathcal{F}_{n})$, $n=N+1,\dots,1$, where
$T_{n}:=E_{n}\dots E_{N},\quad n=N+1,N,\dots,1,$
with $T_{N+1}$ understood to be $1$. (But now we compound $E_{n}$ in a multiplicative manner, while in the previous appendix, e.g., in (32), compounding was additive.) For simplicity, let us assume that all $E_{n}$ are positive, so that $T_{n}$ is a positive martingale.
Let us apply the Shiryaev–Roberts procedure to the e-flow $E_{N},\dots,E_{1},\dots$ continued by setting $E_{0}:=E_{-1}:=\dots:=1$. It gives us the decreasing sequence of stopping times $\sigma_{0}:=N+1$ and
$\sigma_{k}:=\max\left\{n<\sigma_{k-1}\mid\sum_{i=n}^{\sigma_{k-1}-1}E_{n}\dots E_{i}\geq c\right\},\quad k=1,2,\dots,$ (42)
where $n$ ranges over the integers. A useful property of the Shiryaev–Roberts procedure is $\mathbb{E}(\sigma_{k-1}-\sigma_{k}\mid\mathcal{F}_{\sigma_{k}})\geq c$; see, e.g., [28, Proposition 8.13] (proved in [28, Sect. 8.5.2], while the current proof uses other ideas in that section).
To make the stopping times (42) more manageable, it will be convenient to force an alarm every $L$ steps, where $L$ is to be chosen later: $\sigma^{\prime}_{0}:=N+1$ and
$\sigma^{\prime}_{k}:=(\sigma^{\prime}_{k-1}-L)\vee\max\left\{n<\tau_{k-1}\mid\sum_{i=n}^{\sigma^{\prime}_{k-1}-1}E_{n}\dots E_{i}\geq c\right\},\quad k=1,2,\dots;$
let us call this the truncated Shiryaev–Roberts procedure. As in the proof of [28, Corollary 8.4], induction in $k$ shows that $\sigma^{\prime}_{k}\geq\sigma_{k}$ for all $k$.
There is a useful connection between $\tau_{k}$ and $\sigma^{\prime}_{k}$: namely, each set $\{\tau_{k-1}+1,\dots,\tau_{k}\}$ with $\tau_{k}\leq N$ contains at least one stopping time $\sigma^{\prime}_{l}$. (This is even true for $\sigma_{l}$, although we do not need it.) This can be deduced from
$\sum_{j=i}^{\tau_{k}}E_{i}\dots E_{j}\geq c\text{ for some $i\geq\tau_{k-1}+1$}.$ (43)
Indeed, let $l$ be the largest number satisfying $\sigma^{\prime}_{l}>\tau_{k}$; arguing indirectly, let us suppose that $\sigma^{\prime}_{l+1}<\tau_{k-1}+1$. The inequality (43) implies
$\sum_{j=i}^{\sigma^{\prime}_{l}-1}E_{i}\dots E_{j}\geq\sum_{j=i}^{\tau_{k}}E_{i}\dots E_{j}\geq c\text{ for some $i>\sigma^{\prime}_{l+1}$},$
which contradicts the definition of the stopping times $\sigma^{\prime}$.
As in [28, Proof of Proposition 8.15 in Sect. 8.5.2], let us say that $k$ is slow if
$\mathbb{P}\left(\sigma^{\prime}_{k-1}-\sigma^{\prime}_{k}=L\mid\mathcal{F}_{\sigma^{\prime}_{k-1}}\right)\geq c/L$
and fast otherwise. We show there that
$\mathbb{E}\left(\sigma^{\prime}_{k-1}-\sigma^{\prime}_{k}\mid\mathcal{F}_{\sigma^{\prime}_{k-1}}\right)\geq c-c^{2}/L$ (44)
if $k$ is fast. On the other hand, if $k$ is slow,
$\mathbb{E}\left(\sigma^{\prime}_{k-1}-\sigma^{\prime}_{k}\mid\mathcal{F}_{\sigma^{\prime}_{k-1}}\right)\geq L\mathbb{P}\left(\sigma^{\prime}_{k-1}-\sigma^{\prime}_{k}=L\mid\mathcal{F}_{\sigma^{\prime}_{k-1}}\right)\geq c.$
In both cases, we have (44).
Let $A^{\prime}_{N}$ be the largest $k$ such that $\sigma^{\prime}_{k}>0$; we interpret $A^{\prime}_{N}$ as the number of alarms raised by the truncated Shiryaev–Roberts procedure. Remember that $A_{N}$, defined by (28), is the number of alarms raised by the conformal reverse Shiryaev–Roberts e/procedure. By Hoeffding’s inequality [28, Sect. A.6.3], for arbitrarily small $\epsilon\in(0,c)$ and sufficiently large $N$,
$\mathbb{P}\left(\frac{A_{N}}{N}\geq\frac{1}{c-\epsilon}\right)$ $\leq\mathbb{P}\left(\frac{A^{\prime}_{N}}{N}\geq\frac{1}{c-\epsilon}\right)=\mathbb{P}\left(A^{\prime}_{N}\geq\left\lceil\frac{N}{c-\epsilon}\right\rceil\right)$
$\leq\mathbb{P}\left(\sum_{k=1}^{\lceil\frac{N}{c-\epsilon}\rceil}(\sigma^{\prime}_{k-1}-\sigma^{\prime}_{k})\leq N\right)$
$\leq\exp\left(-2\frac{\left(\lceil\frac{N}{c-\epsilon}\rceil\left(c-\frac{c^{2}}{L}\right)-N\right)^{2}}{L^{2}\lceil\frac{N}{c-\epsilon}\rceil}\right)\leq\exp(-\epsilon^{\prime}N),$
where $\epsilon^{\prime}$ is a positive constant (which requires a sufficiently large $L$). Since the series $\sum_{N}\exp(-\epsilon^{\prime}N)$ converges,
$\frac{A_{N}}{N}\geq\frac{1}{c-\epsilon}$
happens only finitely often, which completes the proof of Proposition 19.
To deduce Proposition 18 notice that, by induction in $k$, $\tau_{k}^{\text{SR}}\leq\tau_{k}^{\text{CUSUM}}$, where $\tau_{k}^{\text{CUSUM}}$ are the $\tau_{k}$ defined by (27), and $\tau_{k}^{\text{SR}}$ are the $\tau_{k}$ defined by (30).